OpenStudy (anonymous):

Kendra cuts a triangle out of a piece of paper. With a protractor, she measures one of the angles to be 58.3°. Then she uses a ruler to find that the length of the opposite side is 16.5 cm, and the length of one of the other sides is 18.3 cm. To the nearest tenth of a cm, how long is the third side of Kendra's triangle? A. 13.6 cm B. 18.1 cm C. 15.1 cm D. 18.3 cm please help! I've done all I coud trying to solve this

OpenStudy (anonymous):

looks like a law of sines problem, yes?

OpenStudy (anonymous):



OpenStudy (anonymous):

does the law of sines sound familiar?

OpenStudy (anonymous):

since you posted that here'a another question, how are you suppose to find the third side, and is it the same steps to finding the answer this problem?

OpenStudy (anonymous):

we can find all the angles and all the sides

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{\sin(53.8)}{16.5}=\frac{\sin(B)}{18.3}\] \[\sin(B)=\frac{18.3\sin(53.8)}{16.5}\] \[B=\arcsin(\frac{18.3\sin(53.8)}{16.5})\]

OpenStudy (anonymous):

okay.. I think I see how to do this.

OpenStudy (anonymous):

i get 63.51 http://www.wolframalpha.com/input/?i=arcsin%28 \frac{18.3\sin%2853.8%29}{16.5}%29

OpenStudy (anonymous):

then you know the other angle, since they have to add up to 180 then use the law of sines again to find the remaining side i don't think there is a snappier way to do this, hopefully you are studying law of sines now

OpenStudy (anonymous):

Nope.. That's why for my last two years of high school I am never taking anything geometry related ever again! But I kinda see what's going on here

OpenStudy (anonymous):

who knows, you might end up being a math teacher

OpenStudy (anonymous):

Funny! That's very unlikely..

OpenStudy (anonymous):

not to sound like an old guy, but you never know...

OpenStudy (anonymous):

Whatever happens, happens, hey but thanks for the help man.